alternation voltage V (in volts) is represented by equation V=300 sin(100nt). What is the value of"f" this voltage?


1
Expert's answer
2021-05-04T12:08:00-0400

"\omega=2\pi f \to2\pi f=100\pi\to f=50\ (Hz)"